OpenStudy (theoreo):

Integral of sqrt(1-x^2) I'm not sure if I've learned how to do this and forgotten, or if I never learned how, but I keep getting stuck on integrals involving square roots of polynomials. Can somebody give me a couple of pointers?

OpenStudy (anonymous):

trig sub. 1-x^2 is like a^2 - x^2 so let x=asin(theta) a=1

OpenStudy (theoreo):

Yeah, I was reading through my book and I think I tried it. If \[\sqrt{1-x^2} ; a=1 \] Then \[x = \sin(\theta)\] so \[\sqrt{1-x^2} = \sqrt{1-\sin^2} = \sqrt{\cos^2}\] right?

OpenStudy (anonymous):

yee

OpenStudy (anonymous):

then you just have the integral of cos(theta)

OpenStudy (anonymous):

but the most important thing is to turn your thetas back into x using a right triangle do you know that?

OpenStudy (theoreo):

Which is -sin(theta) right? Ah I forgot the triangle. That's why I wasn't coming up with the right answer.

OpenStudy (anonymous):